OpenStudy (anonymous):

A bread recipe calls for 2/3 cup of white flour and 1 1/2 cups of whole wheat flour. What is the total amount of flour needed. Pick one 1 2/5 cups or 1 3/5 cups or 2 1/6 cups or 1 1/2 cup

OpenStudy (anonymous):

2 1/6 i posted the solution

OpenStudy (anonymous):

2/3 + 3/2 = 13/6 = 2 1/6